# MoDGS: Dynamic Gaussian Splatting from Causually-captured Monocular Videos

### v1

@@ Abstract.
In this paper, we propose MoDGS, a new pipeline to render novel-view images in dynamic scenes using only casually captured monocular videos. Previous monocular dynamic NeRF or Gaussian Splatting methods strongly rely on the rapid movement of input cameras to construct multiview consistency but fail to reconstruct dynamic scenes on casually captured input videos whose cameras are static or move slowly. To address this challenging task, MoDGS adopts recent single-view depth estimation methods to guide the learning of the dynamic scene. Then, a novel 3D-aware initialization method is proposed to learn a reasonable deformation field and a new robust depth loss is proposed to guide the learning of dynamic scene geometry. Comprehensive experiments demonstrate that MoDGS is able to render high-quality novel view images of dynamic scenes from just a casually captured monocular video, which outperforms baseline methods by a significant margin. Project page: https://MoDGS.github.io
Novel View Synthesis, Monocular Video
Figure 1. Given a casually captured monocular video of a dynamic scene, MoDGS is able to synthesize high-quality novel-view images in this scene. Note that in this example, the camera of the input video is static. The baseline methods, i.e. Deformable-GS (Yang et al., 2023a) and SC-GS (Huang et al., 2024), fail to correctly reconstruct the 3D dynamic scenes on this static monocular video. The white regions in cyan bounding boxes are not visible in the input video (red bounding boxes) so there are some artifacts for these invisible regions.
1. Introduction
@@
Novel view synthesis (NVS) is an important task in computer graphics and computer vision, which greatly facilitates downstream tasks such as augmentation or virtual reality. In recent years, the novel-view-synthesis quality on static scenes has witnessed great improvements thanks to the recent development of techniques such as NeRF (Mildenhall et al., 2020), Instant-NGP (Müller et al., 2022), and Gaussian Splatting (Kerbl et al., 2023), especially when there are sufficient input images. However, novel view synthesis in a dynamic scene with only one monocular video still remains a challenging task.
@@
Dynamic View Synthesis (DVS) has achieved impressive improvements along with the emerging neural representations and Gaussian splitting techniques. Most of the existing DVS methods (Cao and Johnson, 2023; Yang et al., 2023a) require multiview videos captured by dense synchronized cameras to achieve good rendering quality. Though some works can process a monocular video for DVS, as pointed out by DyCheck (Gao et al., 2022), these methods require the camera of the monocular video to have extremely large movements, which is called ”Teleporting Camera Motion” on different viewpoints, so these methods can utilize the multiview consistency provided by this pseudo multiview video to reconstruct the 3D geometry of the dynamic scene. However, such large camera movements are rarely seen in casually captured videos because casually captured videos are usually produced by smoothly moving or even static cameras. When the camera moves slowly or is static, the multiview consistency constraint will be much weaker and all these existing DVS methods fail to produce high-quality novel-view images, as shown in Fig. 1 (second and third column).
@@
In this paper, we present Monocular Dynamic Gaussian Splatting (MoDGS) to render novel-view images from casually captured monocular videos in a dynamic scene. MoDGS addresses the weak multiview constraint problem by adopting a monocular depth estimation method  (Fu et al., 2024). The weak multiview constraint problem disables existing methods to correctly reconstruct the 3D dynamic scenes while the single view depth estimator provides prior depth information on the input video to help the 3D reconstruction. However, we find that simply applying a single-view depth estimator in DVS to supervise rendered depth maps is not enough for high-quality novel view synthesis. First, the depth supervision only provides information for each frame but does not help to associate 3D points between two frames. Thus, we still have difficulty in learning an accurate time-dependent deformation field. Second, the estimated depth values are not consistent among different frames.
@@
To learn a robust deformation field from monocular video, we propose a 3D-aware initialization scheme for the deformation field. Existing methods (Katsumata et al., 2023) solely rely on supervision from 2D flow estimation, which produces deteriorated results without sufficient multiview consistency. We find that directly initializing the deformation field in the 3D space greatly helps the subsequent learning of the 3D representations and improves the rendering quality in the end(Fig. 1,last column).
@@
To better utilize the estimated depth maps for supervision, we propose a novel depth loss to address the scale inconsistency of estimated depth values across different frames. Previous methods (Li et al., 2023b; Liu et al., 2023a) supervise the rendered depth maps using a scale-invariant depth loss by minimizing the $L2$ distance of normalized rendered depth and depth priors, and the most recent method  (Zhu et al., 2023)propose to supervise the rendered depth maps using a Pearson correlation loss to mitigate the scale ambiguity between the true scene scale and the estimate depth scale. However, the estimated depth maps of different frames are not even consistent after normalizing to the same scale. We observe that despite the inconsistency in values, the orders of depth values of different pixels in different frames are stable, which motivates us to propose an ordinal depth loss. This novel ordinal depth loss enables us to fully utilize the estimated depth maps for high-quality novel view synthesis.
@@
To demonstrate the effectiveness of MoDGS, we conduct experiments on two widely used datasets, the Nvdia (Yoon et al., 2020) dataset and the DyNeRF (Li et al., 2022) dataset, and a self-collected dataset containing monocular videos from the Internet. We adopt an exact monocular DVS evaluation setting that only uses the video of one camera as input while evaluating the video of another camera. Results show that our method outperforms previous DVS methods by a large margin and achieves high-quality NVS on casually captured monocular videos.
Figure 2. Overview. Given a casually captured monocular video of a dynamic scene, MoDGS represents the dynamic scene with a set of Gaussians in a canonical space and a deformation field represented by an MLP network $\mathcal{T}$. To render an image at a specific timestep $t$, we deform all the Gaussians by $\mathcal{T}_{t}$ and then use the splatting technique to render images and depth maps. While in training MoDGS, we use a single-view depth estimator GeoWizard (Fu et al., 2024) to estimate a depth map for every frame and compute the rendering loss and an ordinal depth loss to learn MoDGS.
2. Related works
@@
2.1. Dynamic NeRF
@@
Recent dynamic NeRF methods can be roughly categorized into two groups. 1) Representing by time-varying neural radiance fields conditioned on time (Gao et al., 2021; Li et al., 2022; Park et al., 2023). For example, Park et al. (2023) proposes a simple spatiotemporal radiance field by interpolating the feature vectors indexed by time. 2) Representing by a canonical space NeRF and deformation field (Guo et al., 2023; Li et al., 2021; Park et al., 2021a, b; Pumarola et al., 2021; Tretschk et al., 2021; Xian et al., 2021). For example, NSFF (Li et al., 2021) models the dynamic components using forward and backward flow represented as 3D dense vector fields; Nerfies (Park et al., 2021a) and HyperNeRF (Park et al., 2021b) model the scene dynamics as a deformation field mapping to a canonical space. Recent advances in grid-based NeRFs (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022; Chen et al., 2022) demonstrate that the training of static NeRFs can be significantly accelerated. Consequently, some dynamic NeRF works utilize these grid-based or hybrid representations for fast optimization (Guo et al., 2023; Cao and Johnson, 2023; Fang et al., 2022; Fridovich-Keil et al., 2023; Shao et al., 2023; Wang et al., 2023c, b; Song et al., 2023; You and Hou, 2023).
2.2. Dynamic Gaussian Splatting
@@
The recent emergence of 3D Gaussian Splatting (3DGS) demonstrates its efficacy for super-fast real-time rendering attributed to its explicit point cloud representation. Recent follow-ups extend 3DGS to model dynamic 3D scenes. Luiten et al. (2023) and Katsumata et al. (2023) track dynamic 3D Gaussians by frame-by-frame training from synchronized multi-view videos. Yang et al. (2023a) propose a deformable version of 3DGS by introducing a deformation MLP network to model the 3D flows. Wu et al. (2023) and Duisterhof et al. (2023) also introduce a deformation field but using a more efficient Hexplane representation (Cao and Johnson, 2023). Yang et al. (2023b) proposes a dynamic representation with a collection of 4D Gaussian primitives, where the time evolution can be encoded by 4D spherical harmonics. Bae et al. (2024) encodes motions with a per-Gaussian feature vector. Some other works (Li et al., 2023a; Lin et al., 2023; Liang et al., 2023) also study how to effectively encode the motions for Gaussians with different bases. To effectively learn the motions of Gaussians, some works (Feng et al., 2023; Yu et al., 2023; Huang et al., 2024) resort to clustering the motions together for a compact representation.
2.3. DVS from Casual Monocular Videos
@@
As presented in DyCheck (Gao et al., 2022), most existing monocular dynamic view synthesis datasets commonly used for benchmarking, such as D-NeRF (Pumarola et al., 2021), HyperNeRF (Park et al., 2021b), and Nerfies (Park et al., 2021a), either ”teleport” frames between multiple cameras (i.e., involve large camera movements between neighboring frames) or feature quasi-static object motions (i.e., small object motions over a long duration). This frame-capturing style facilitates the utilization of multi-view constraints, enabling easy dynamic 3D modeling; however, it is not common in daily casual video captures. When taking casual videos as input, quality degradations appear in the reconstruction results produced by the aforementioned works. A few works explore robust dynamic 3D scene modeling given monocular, casual videos. DynIBaR (Li et al., 2023b) enables long-sequence image-based rendering of dynamic scenes by aggregating features from nearby views following a scene motion-aware manner, but the training cost is large for long-time per-scene optimization. Lee et al. (2023) proposes a hybrid representation that combines static and dynamic components, enabling both fast training and rendering. However, it requires per-frame masks indicating the dynamic components as an additional input. RoDynRF (Liu et al., 2023a) enables robust dynamic NeRF reconstruction by jointly estimating NeRF parameters and camera pose parameters. DpDy (Wang et al., 2024) introduces additional supervision by fine-tuning a diffusion model and imposing this supervision using the SDS loss (Poole et al., 2022), where the quality heavily relies on diffusion models, which require a large amount of computational resources. In contrast, our method is more lightweight, as it does not introduce large neural networks and offers better computing efficiency.
@@
Figure 3. (a) Initialization of the deformation field. We first lift the depth maps and a 2D flow to a 3D flow and train the deformation field for initialization. (b) Initialization of Gaussians in the canonical space. We use the initialized deformation field to deform all the depth points to the canonical space and downsample these depth points to initialize Gaussians.
3. Methodology
@@
Given a casually captured monocular video, we aim to synthesize novel view images from this video. We propose MoDGS, which achieves this by learning a set of Gaussians $\{G_{i}|i=1,~{}\cdots,~{}N\}$ in a canonical space and a deformation field $\mathcal{T}_{t}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ to deform these Gaussians to a specific timestamp $t$. Then, for a specific timestamp $t$ and a camera pose, we will adopt the splitting technique to render an image.
Overview
@@
As shown in Fig. 2, to train MoDGS, we split the monocular video into a sequence of images $\{I_{t}|t=1,...,T\}$ and we assume that all camera poses of all images are known. We denote our deformation field as a function $x_{t}=\mathcal{T}_{t}(x)$, which maps a 3D location $x\in\mathbb{R}^{3}$ in the canonical 3D space to a location $x_{t}\in\mathbb{R}^{3}$ in the 3D space on time $t$. For every image $I_{t}$, we utilize a single-view depth estimator GeoWizard (Fu et al., 2024) to estimate a depth map $D_{t}$ for every image and utilize a flow estimation method RAFT (Teed and Deng, 2020) to estimate a 2D optical flow $F_{t_{i}\to t_{j}}$ between $I_{t_{i}}$ and $I_{t_{j}}$ where $t_{i}$ and $t_{j}$ are two arbitrary timesteps. Then, we initialize our deformation field by a 3D-aware initialization scheme as introduced in Sec. 3.2. After initialization, we train our Gaussians and deformation field with a rendering loss and a new depth loss introduced in Sec. 3.3. In the following, we first begin with the definition of the Gaussians and the rendering process in MoDGS.
3.1. Gaussians and Deformation Fields
Gaussians in the canonical space

### v2

# MoDGS: Dynamic Gaussian Splatting from Casually-captured Monocular Videos

@@ Abstract
In this paper, we propose MoDGS, a new pipeline to render novel views of dynamic scenes from a casually captured monocular video. Previous monocular dynamic NeRF or Gaussian Splatting methods strongly rely on the rapid movement of input cameras to construct multiview consistency but struggle to reconstruct dynamic scenes on casually captured input videos whose cameras are either static or move slowly. To address this challenging task, MoDGS adopts recent single-view depth estimation methods to guide the learning of the dynamic scene. Then, a novel 3D-aware initialization method is proposed to learn a reasonable deformation field and a new robust depth loss is proposed to guide the learning of dynamic scene geometry. Comprehensive experiments demonstrate that MoDGS is able to render high-quality novel view images of dynamic scenes from just a casually captured monocular video, which outperforms state-of-the-art methods by a significant margin. The code will be publicly available. Project page:   https://MoDGS.github.io
1 Introduction
@@
Novel view synthesis (NVS) is an important task in computer graphics and computer vision, which greatly facilitates downstream applications such as augmented or virtual reality. In recent years, the novel-view-synthesis quality on static scenes has witnessed great improvements thanks to the recent development of techniques such as NeRF (Mildenhall et al., 2020), Instant-NGP (Müller et al., 2022), and Gaussian Splatting (Kerbl et al., 2023), especially when there are sufficient input images. However, novel view synthesis in a dynamic scene with only one monocular video still remains a challenging task.
@@
Dynamic View Synthesis (DVS) has achieved impressive improvements along with the emerging neural representations (Mildenhall et al., 2020) and Gaussian splatting (Kerbl et al., 2023) techniques. Most of the existing DVS methods (Cao & Johnson, 2023; Yang et al., 2023b) require multiview videos captured by dense synchronized cameras to achieve good rendering quality. Though some works can process a monocular video for DVS, as pointed out by DyCheck (Gao et al., 2022), these methods require the camera of the monocular video to have extremely large movements, which is called “Teleporting Camera Motion” on different viewpoints, so these methods can utilize the multiview consistency provided by this pseudo multiview video to reconstruct the 3D geometry of the dynamic scene. However, such large camera movements are rarely seen in casually captured videos because casual videos are usually produced by smoothly moving or even static cameras. When the camera moves slowly or is static, the multiview consistency constraint will be much weaker and all these existing DVS methods fail to produce high-quality novel-view images, as shown in Fig. 1.
@@
In this paper, we present Monocular Dynamic Gaussian Splatting (MoDGS) to render novel-view images from casually captured monocular videos in a dynamic scene. MoDGS addresses the weak multiview constraint problem by adopting a monocular depth estimation method  (Fu et al., 2024), which provides prior depth information on the input video to help the 3D reconstruction.
However, we find that simply applying a single-view depth estimator in DVS to supervise rendered depth maps is not enough for high-quality novel view synthesis. First, the depth supervision only provides information for each frame but does not help to associate 3D points between two frames in time. Thus, we still have difficulty in learning an accurate time-dependent deformation field. Second, the estimated depth values are not consistent among different frames.
@@
To learn a robust deformation field from a monocular video, we propose a 3D-aware initialization scheme for the deformation field. Existing methods (Katsumata et al., 2023) solely rely on supervision from 2D flow estimation, which produces deteriorated results without sufficient multiview consistency. We find that directly initializing the deformation field in the 3D space greatly helps the subsequent learning of the 4D representations and improves the rendering quality as shown in Fig. 1.
To better utilize the estimated depth maps for supervision, we propose a novel depth loss to address the scale inconsistency of estimated depth values across different frames.
@@
Previous methods (Li et al., 2023b; Liu et al., 2023a) supervise the rendered depth maps using a scale-invariant depth loss by minimizing the $L2$ distance of normalized rendered depth and depth priors, and the most recent method  (Zhu et al., 2023) proposed to supervise the rendered depth maps using a Pearson correlation loss to mitigate the scale ambiguity between the reconstructed scene and the estimated depth maps. However, the estimated depth maps of different frames are not even consistent after normalizing to the same scale. To address these challenges, we observe that despite the inconsistency in values, the orders of depth values of different pixels in different frames are stable, which motivates us to propose an ordinal depth loss. This novel ordinal depth loss enables us to fully utilize the estimated depth maps for high-quality novel view synthesis.
@@
To demonstrate the effectiveness of MoDGS, we conduct experiments on three widely used datasets, the Nvdia (Yoon et al., 2020) dataset, the DyNeRF (Li et al., 2022) dataset, and the Davis (Pont-Tuset et al., 2017) dataset. We also present results on a self-collected dataset containing monocular in-the-wild videos from the Internet. We adopt an exact monocular DVS evaluation setting that only uses the video of one camera as input while evaluating the video of another camera. Results show that our method outperforms previous DVS methods by a large margin and achieves high-quality NVS on casually captured monocular videos.
Figure 1: Given a casually captured monocular video of a dynamic scene, MoDGS is able to synthesize high-quality novel-view images in this scene. In the middle column, the baseline method (Yang et al., 2023b) fails to correctly reconstruct the 3D dynamic scenes on this static monocular video. The white regions in cyan bounding boxes are not visible in the input video (red bounding boxes) so there are some artifacts for these invisible regions. In the rightmost column, the input estimated monocular depth is inconsistent (red bounding boxes); however, our proposed ordinal depth loss effectively ensures more consistent depth outputs. This loss enhances the accuracy and reliability of learning underlying geometry.
2 Related Work
@@
Figure 2: Overview. Given a casually captured monocular video of a dynamic scene, MoDGS represents the dynamic scene with a set of Gaussians in a canonical space and a deformation field represented by an MLP $\mathcal{T}$. To render an image at a specific timestamp $t$, we deform all the Gaussians by $\mathcal{T}_{t}$ and then use the splatting technique to render images and depth maps. While in training MoDGS, we use a single-view depth estimator GeoWizard (Fu et al., 2024) to estimate depth maps and compute the rendering loss and an ordinal depth loss for training.
@@
Dynamic NeRF. Recent dynamic NeRF methods can be roughly categorized into two groups. 1) Representing by time-varying neural radiance fields conditioned on time (Gao et al., 2021; Li et al., 2022; Park et al., 2023). For example, Park et al. (2023) proposes a simple spatiotemporal radiance field by interpolating the feature vectors indexed by time. 2) Representing by a canonical space NeRF and deformation field (Guo et al., 2023; Li et al., 2021; Park et al., 2021a; b; Pumarola et al., 2021; Tretschk et al., 2021; Xian et al., 2021). For example, NSFF (Li et al., 2021) models the dynamic components using forward and backward flow represented as 3D dense vector fields; Nerfies (Park et al., 2021a) and HyperNeRF (Park et al., 2021b) model the scene dynamics as a deformation field mapping to a canonical space. Recent advances in grid-based NeRFs (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022; Chen et al., 2022) demonstrate that the training of static NeRFs can be significantly accelerated. Consequently, some dynamic NeRF works utilize these grid-based or hybrid representations for fast optimization (Guo et al., 2023; Cao & Johnson, 2023; Fang et al., 2022; Fridovich-Keil et al., 2023; Shao et al., 2023; Wang et al., 2023a; b; Song et al., 2023; You & Hou, 2023).
@@
Dynamic Gaussian Splatting. The recent emergence of 3D Gaussian Splatting (3DGS) demonstrates its efficacy for super-fast real-time rendering attributed to its explicit point cloud representation. Recent follow-ups extend 3DGS to model dynamic 3D scenes. Luiten et al. (2023) track dynamic 3D Gaussians by frame-by-frame training from synchronized multi-view videos. Yang et al. (2023b) propose a deformable version of 3DGS by introducing a deformation MLP network to model the 3D flows. Wu et al. (2023) and Duisterhof et al. (2023) also introduce a deformation field but using a more efficient Hexplane representation (Cao & Johnson, 2023). Yang et al. (2023a) proposes a dynamic representation with a collection of 4D Gaussian primitives, where the time evolution can be encoded by 4D spherical harmonics. Bae et al. (2024) encodes motions with a per-Gaussian feature vector. Some other works (Li et al., 2023a; Lin et al., 2023; Liang et al., 2023) also study how to effectively encode the motions for Gaussians with different bases. To effectively learn the motions of Gaussians, some works (Feng et al., 2023; Yu et al., 2023; Huang et al., 2024) resort to clustering the motions together for a compact representation.
@@
DVS from Casual Monocular Videos. As shown in DyCheck (Gao et al., 2022), many existing monocular dynamic view synthesis datasets used for benchmarking, like D-NeRF (Pumarola et al., 2021), HyperNeRF (Park et al., 2021b), and Nerfies (Park et al., 2021a), typically involve significant camera movements between frames but with small object dynamic motions. While this capture style helps with multi-view constraints and dynamic 3D modeling, it is not representative of casual everyday video captures. When using casual videos, the reconstruction results from these methods suffer from quality degradation. Some works address dynamic 3D scene modeling using monocular casual videos. DynIBaR (Li et al., 2023b) allows for long-sequence image-based rendering of dynamic scenes by aggregating features from nearby views, but its training cost is high for long per-scene optimization. Lee et al. (2023) proposes a hybrid representation that combines static and dynamic elements, allowing for faster training and rendering, though it requires additional per-frame masks for dynamic components. RoDynRF (Liu et al., 2023a) focuses on robust dynamic NeRF reconstruction by estimating NeRF and camera parameters together. DpDy (Wang et al., 2024a) enhances quality by fine-tuning a diffusion model with SDS loss supervision (Poole et al., 2022), but it demands significant computational resources. Concurrent works like DG-Marbles (Stearns et al., 2024) use Gaussian marbles and a hierarchical learning strategy to optimize representations. Shape-of-Motion (Wang et al., 2024b) and Mosca (Lei et al., 2024) rely on explicit motion representation and initialize scene deformation with depth estimation and video tracking priors. However, our MoDGS method effectively uses only noisy inter-frame flow maps from RAFT (Teed & Deng, 2020) as input, performing well without the need for strong long-range pixel correspondence.
Ordinal Relation in Depth Maps.
The ordinal relation between pixels has been investigated in recent years, especially in the field of monocular depth estimation.  Zoran et al. (2015) proposed to use a three-category classification network to predict the order relation of given pixel pairs. Then the depth can be extracted by optimizing a constrained quadratic optimization problem. Similarly,  Fu et al. (2018) treats the depth prediction problem as a multi-class classification problem.
Chen et al. (2016) further proposes a ranking loss to learn metric depth, which encourages a small difference between depths if the ground-truth relation is equal; otherwise it encourages a large difference. Then, Pavlakos et al. (2018) extends this differentiable ranking loss to the human pose estimation task. However, these works only utilize limited numbers of depth orders for training (one pair in  Chen et al. (2016) and 17 pairs in  Pavlakos et al. (2018)), resulting in coarse supervision for depth maps. The direct application of their ranking loss as depth supervision has yet to be explored. Moreover, our ordinal depth takes the rendered metric depth maps as input, which are dense grids of float numbers. Our task is different from previous depth estimation and pose estimation tasks and we present a comparison between our ordinal depth loss and depth ranking loss in Appendix A.10.
@@
Figure 3: (a) Initialization of the deformation field. We first lift the depth maps and a 2D flow to a 3D flow and train the deformation field for initialization. (b) Initialization of Gaussians in the canonical space. We use the initialized deformation field to deform all the depth points to the canonical space and downsample these depth points to initialize Gaussians.
3 Proposed Method
@@
Given a casually captured monocular video, we aim to synthesize novel view images from this video. We propose MoDGS, which achieves this by learning a set of Gaussians $\{G_{i}|i=1,~{}\cdots,~{}N\}$ in a canonical space and a deformation field $\mathcal{T}_{t}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ to deform these Gaussians to a specific timestamp $t$. Then, for a timestamp $t$ and a camera pose, we use splatting to render an image.
Overview.
@@
As shown in Fig. 2, to train MoDGS, we split the monocular video into a sequence of images $\{I_{t}|t=1,...,T\}$ with known camera poses. We denote our deformation field as a function $x_{t}=\mathcal{T}_{t}(x)$, which maps a 3D location $x\in\mathbb{R}^{3}$ in the canonical 3D space to a location $x_{t}\in\mathbb{R}^{3}$ in the 3D space on time $t$. For every image $I_{t}$, we utilize a single-view depth estimator (Fu et al., 2024) to estimate a depth map $D_{t}$ for every image and utilize a flow estimation method RAFT (Teed & Deng, 2020) to estimate a 2D optical flow $F_{t_{i}\to t_{j}}$ between $I_{t_{i}}$ and $I_{t_{j}}$ where $t_{i}$ and $t_{j}$ are two arbitrary timestamps. Then, we initialize our deformation field by a 3D-aware initialization scheme as introduced in Sec. 3.2. After initialization, we train our Gaussians and deformation field with a rendering loss and a new depth loss introduced in Sec. 3.3. In the following, we first begin with the definition of the Gaussians and the rendering process in MoDGS.
3.1 Gaussians and Deformation Fields
Gaussians in the canonical space.